According to a report by the Commerce Department in the fall of 2004, 20% of U.S. households had some type of high-speed Internet connection. Let Nn denote the number of U.S. households with a high-speed Internet connection in n households. What is the probability that more than 10 of the first 100 households selected have high-speed Internet?AA
Answer #1

This is the binomial problem

n = 100 , p = 0.20
As per binomial distribution,
P(X=r) = nCr * p^r * (1-p)^(n-r)

P(x>10) = 1- P(x< 10)
= 1- 0.0057
= 0.9943